# MsgABD: EveryShipout initializing macros

### v1

@@ A.1 Methods for Alleviating Over-smoothing
A popular choice for reducing over-smoothing in GNNs is to regularize the model. Recall that DropEdge (Rong et al., 2020) implicitly regularizes the model by adding noise to the learning trajectory (Section 2.2). Similarly, Dropout Srivastava et al. (2014) drops entries from the input features to each layer, and DropMessage (Fang et al., 2023) drops entries from the message matrix (before the aggregation step). Graph Drop Connect (GDC) (Hasanzadeh et al., 2020) combines DropEdge and DropMessage together, resulting in a layer-wise sampling scheme that uses a different subgraph for message-aggregation over each feature dimension. These methods successfully addressed the over-smoothing problem, enabling the training of deep GNNs, and performed competitively on several benchmarking datasets.
@@
Another powerful form of implicit regularization is feature normalization, which has proven crucial in enhancing the performance and stability of several types of neural networks (Huang et al., 2020). Exploiting the inductive bias in graph-structured data, normalization techniques like PairNorm (Zhao & Akoglu, 2020), Differentiable Group Normalization (DGN) (Zhou et al., 2020) and NodeNorm (Zhou et al., 2021b) have been proposed to reduce over-smoothing in GNNs. On the other hand, Energetic Graph Neural Networks (EGNNs) Zhou et al. (2021a) explicitly regularize the optimization by constraining the layer-wise Dirichlet energy to a predefined range.
@@
In a different vein, motivated by the success of residual networks (ResNets) (He et al., 2016) in computer vision, Li et al. (2019) proposed the use of residual connections to prevent the smoothing of representations. Residual connections successfully improved the performance of GCN on a range of graph-learning tasks. Chen et al. (2020) introduced GCNII, which uses skip connections from the input to all hidden layers. This layer wise propagation rule has allowed for training of ultra-deep networks – up to 64 layers. Some other architectures, like the Jumping Knowledge Network (JKNet) Xu et al. (2018) and the Deep Adaptive GNN (DAGNN) Liu et al. (2020), aggregate the representations from all layers, $\{\bm{z}^{\left(\ell\right)}_{i}\}_{\ell=1}^{L}$, before processing them through a readout layer.
A.2 Methods for Alleviating Over-squashing
@@
Spatial rewiring methods use the topological relationships between the nodes in order to come up with a rewiring strategy. That is the graph rewiring is guided by the objective of optimizing some chosen topological properties. For instance, Alon & Yahav (2021) introduced a fully-adjacent (FA) layer, wherein messages are passed between all nodes. GNNs using a FA layer in the final message-passing step were shown to outperform the baselines on a variety of long-range tasks, revealing the importance of information exchange between far-off nodes which standard message-passing cannot facilitate. Topping et al. (2022) proposed a curvature-based rewiring strategy, called the Stochastic Discrete Ricci Flow (SDRF), which aims to reduce the “bottleneckedness” of a graph by adding suitable edges, while simultaneously removing edges in an effort to preserve the statistical properties of the original topology. Black et al. (2023) proposed the Greedy Total Resistance (GTR) technique, which optimizes the graph’s total resistance by greedily adding edges to achieve the greatest improvement. One concern with graph rewiring methods is that unmoderated densification of the graph, e.g. using a fully connected graph for propagating messages, can result in a loss of the inductive bias the topology provides, potentially leading to over-fitting. Accordingly, Gutteridge et al. (2023) propose a Dynamically Rewired (DRew) message-passing framework that gradually densifies the graph. Specifically, in a given layer $\ell$, node $i$ aggregates messages from its entire $\ell$-hop receptive field instead of just the immediate neighbors. This results in an improved communication over long distances while also retaining the inductive bias of the shortest distance between nodes.
@@
Some other rewiring techniques don’t exactly classify as spatial or spectral methods. For instance, Probabilistically Rewired MPNN (PR-MPNN) Qian et al. (2024) learns to probabilistically rewire a graph, effectively mitigating under-reaching as well as over-squashing. Finally, Gabrielsson et al. (2023) proposed connecting all nodes at most $r$-hops away, for some $r\in\mathbb{N}$, and introducing positional embeddings to allow for distance-aware aggregation of messages.
A.3 Towards a Unified Treatment
@@
Several studies have shown that an inevitable trade-off exists between the problems of over-smoothing and over-squashing, meaning that optimizing for one will compromise the other. For instance, Topping et al. (2022); Nguyen et al. (2023) showed that negatively curved edges create bottlenecks in the graph resulting in over-squashing of information. On the other hand, Nguyen et al. (2023, Proposition 4.3) showed that positively curved edges in a graph contribute towards the over-smoothing problem. To address this trade-off, they proposed Batch Ollivier-Ricci Flow (BORF), which adds new edges adjacent to the negatively curved ones, and simultaneously removes positively curved ones. In a similar vein, Giraldo et al. (2023) demonstrated that the minimum number of message-passing steps required to reach a given level of over-smoothing is inversely related to the Cheeger’s constant, $h_{\mathcal{G}}$. This again implies an inverse relationship between over-smoothing and over-squashing. To effectively alleviate the two issues together, they proposed the Stochastic Jost and Liu Curvature Rewiring (SJLR) algorithm, which adds edges that result in high improvement in the curvature of existing edges, while simultaneously removing those that have low curvature.
@@
Despite the well-established trade-off between over-smoothing and over-squashing, some works have successfully tackled them together despite only adding or removing edges. One such work is Karhadkar et al. (2023), which proposed a rewiring algorithm that adds edges to the graph but does not remove any. The First-order Spectral Rewiring (FoSR) algorithm computes, as the name suggests, a first order approximation to the spectral gap of $\bm{L}^{\text{sym}}$, and adds edges with the aim of maximizing it. Since the spectral gap directly relates to Cheeger’s constant through Cheeger’s inequality (Alon & Milman, 1985; Alon, 1986; Sinclair & Jerrum, 1989), this directly decreases the over-squashing levels. Moreover, Karhadkar et al. (2023, Figure 5) empirically demonstrated that addition of (up to a small number of) edges selected by FoSR can lower the Dirichlet energy of the representations. Taking a completely opposite approach, CurvDrop (Liu et al., 2023) adapted DropEdge to remove negatively curved edges sampled from a distribution proportional to their curvatures. CurvDrop directly reduces over-squashing and, as a side benefit of operating on a sparser subgraph, also mitigates over-smoothing.
@@ Appendix B Proofs
B.1 Expected Propagation Matrix under DropEdge
@@ B.2 Sensitivity in L-layer Linear GCNs
where $\mathsf{ND}$ refers to a NoDrop model ($q=0$), the placeholder $\cdots$ can be replaced with one of the edge-dropping methods $\mathsf{DE}$, $\mathsf{DA}$ or $\mathsf{DG}$, and the corresponding entries of $\mathbb{E}_{\ldots}[\hat{\bm{A}}]$ can be plugged in from Equation 3.3, Equation 3.5 and Equation 3.6, respectively. Moreover, the sensitivity monotonically decreases as the dropping probability is increased.
@@
$\mathsf{Paths}\left(j\to i\right)=\left\{\left(u_{0},\ldots,u_{L}\right):u_{0}=j;u_{L}=i;\left(u_{\ell-1}\to u_{\ell}\right)\in\mathcal{E},\forall\ell\in\left[L\right]\right\}$
@@
The right hand side of the inequality is the $\left(i,j\right)$-entry in the Lth power of the propagation matrix of a NoDrop model. From Equation 3.5 and Equation 3.6, we know that Equation B.17 is true for DropAgg and DropGNN as well. We conclude the first part of the proof using Equation 3.2 – the sensitivity of node $i$ to node $j$ is proportional to $(\dot{\bm{P}}^{L})_{ij}$.
@@
Each of the terms on the right are increasing in $q$, hence, all the $\dot{\bm{P}}_{u_{\ell}u_{\ell-1}}$ factors are decreasing in $q$. Using this result with Equation B.15, we conclude the second part of the theorem.
Appendix C Test-time Monte-Carlo Averaging
In Section 3, we focused on the expected sensitivity of the stochastic representations in models using DropEdge-like regularization strategies. This corresponds to their training-time behavior, wherein the activations are random. At test-time, the standard practice is to turn these methods off by setting $q=0$. However, this raises the over-smoothing levels back up Xuanyuan et al. (2023). Another way of making predictions is to perform multiple stochastic forward passes, as during training, and then averaging the model outputs. This is similar to Monte-Carlo Dropout, which is an efficient way of ensemble averaging in MLPs Gal & Ghahramani (2016b), CNNs Gal & Ghahramani (2016a) and RNNs Gal & Ghahramani (2016c). In addition to alleviating over-smoothing, this approach also outperforms the standard implementation Xuanyuan et al. (2023). We can study the effect of random edge-dropping in this setting by examining the sensitivity of the expected representations:
@@ Appendix C Test-time Monte-Carlo Averaging
$\left\|\frac{\partial}{\partial\bm{x}_{j}}\mathbb{E}\left[\bm{z}^{\left(L\right)}_{i}\right]\right\|_{1}$
In linear models, the order of the two operations – expectation and sensitivity computation – is irrelevant:
@@
$\mathbb{E}\left[\left\|\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right\|_{1}\right]=\left\|\mathbb{E}\left[\hat{\bm{A}}_{ij}\right]\bm{W}\right\|_{1}=\left\|\mathbb{E}\left[\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right]\right\|_{1}=\left\|\frac{\partial}{\partial\bm{x}_{j}}\mathbb{E}\left[\bm{z}^{\left(L\right)}_{i}\right]\right\|_{1}$ (C.1)

### v2

@@ A.1 Methods for Alleviating Over-smoothing
A popular choice for reducing over-smoothing in GNNs is to regularize the model. Recall that DropEdge [68] implicitly regularizes the model by adding noise to it (Section 2.2). A similarly regularization effect is observed with the methods discussed in the main text – DropNode [23], DropAgg [43], DropGNN [65], Dropout [77] and DropMessage [21]. Graph Drop Connect (GDC) [37] combines DropEdge and DropMessage together, resulting in a layer-wise sampling scheme that uses a different subgraph for message-aggregation over each feature dimension. These methods successfully addressed the over-smoothing problem, enabling the training of deep GNNs, and performed competitively on several benchmarking datasets.
@@
Another powerful form of implicit regularization is feature normalization, which has proven crucial in enhancing the performance and stability of several types of neural networks [41]. Exploiting the inductive bias in graph-structured data, normalization techniques like PairNorm [93], Differentiable Group Normalization (DGN) [95] and NodeNorm [97] have been proposed to reduce over-smoothing in GNNs. On the other hand, Energetic Graph Neural Networks (EGNNs) [96] explicitly regularize the optimization by constraining the layer-wise Dirichlet energy to a predefined range.
@@
In a different vein, motivated by the success of residual networks (ResNets) [38] in computer vision, Li et al. [49] proposed the use of residual connections to prevent the smoothing of representations. Residual connections successfully improved the performance of GCN on a range of graph-learning tasks. Chen et al. [13] introduced GCN-II, which uses skip connections from the input to all hidden layers. This layer wise propagation rule has allowed for training of ultra-deep networks – up to 64 layers. Some other architectures, like the Jumping Knowledge Network (JKNet) [84] and the Deep Adaptive GNN (DAGNN) [53], aggregate the representations from all layers, $\{\bm{z}^{\left(\ell\right)}_{i}\}_{\ell=1}^{L}$, before processing them through a readout layer.
A.2 Homophily Bias in Evaluation of Techniques for Deep GNN
Table 3: Statistics of node-classification datasets. Homophily measures as defined in Lim et al. [52].
Dataset Nodes Edges Features Classes Homophily
Homophilic Networks
Reddit 232,965 114,615,892 602 41 0.653
OGBN-ArXiv 169,343 1,166,243 128 40 0.416
Coauthor-CS 18,333 163,788 6,805 15 0.755
Coauthor-Physics 34,493 495,924 8,415 5 0.847
Wiki-CS 11,701 216,123 300 10 0.568
Amazon-Computers 13,752 491,722 767 10 0.700
Amazon-Photo 7,650 238,162 745 8 0.772
Heterophilic Networks
Flickr 89,250 899,756 500 7 0.070
Cornell 183 298 1,703 5 0.031
Texas 183 325 1,703 5 0.001
Wisconsin 251 515 1,703 5 0.094
We examine the evaluation protocols commonly used for assessing methods aimed at alleviating over-smoothing in deep GNNs – many of which are also widely adopted for training deep architectures. Notably, we highlight a misalignment between the intended goal of these methods – to improve the trainability of deep GNNs – and their evaluation, which is often restricted to short-range tasks.
For example, DropEdge [68] was evaluated on Cora [57], CiteSeer [31], PubMed [62], and a version of Reddit [36] distinct from the one used in our experiments. The first three exhibit high label homophily (see Table 4) and are known to be better modelled by shallower networks [95]. Reddit also displays strong homophily, as can be seen in Table 3. Similarly, DropNode [23] was evaluated on Cora, CiteSeer, and PubMed; DropAgg [43] on Cora ML, CiteSeer, and OGBN-ArXiv [40], which has moderate homophily; DropMessage [68] was evaluated on Cora, CiteSeer, PubMed, OGBN-ArXiv, and Flickr, with only the latter having low homophily; GDC [37] was evaluated on Cora, Cora ML and CiteSeer.
A similar trend can be observed in the evaluation of feature normalization techniques used to regularize GNNs. PairNorm [93] and DGN [95] were evaluated on Cora, CiteSeer, PubMed, and Coauthor-CS [75]; NodeNorm [97] on Cora, CiteSeer, PubMed, Coauthor-CS, Wiki-CS [58], and Amazon-Photo [75]; and EGNNs [96] on Cora, PubMed, Coauthor-Physics [75], and OGBN-ArXiv – all of these datasets are highly homophilic.
A similar trend is observed in the evaluation of architectural modifications designed to enable deeper GNNs. GCN-II [13] on Cora, CiteSeer, PubMed, and Chameleon; JKNet [84] on Cora, CiteSeer, and Reddit; and DAGNN [53] on Cora, CiteSeer, PubMed, Coauthor-CS, Coauthor-Physics, Amazon-Computers [75], and Amazon-Photo – many of these datasets are highly homophilic as well.
This pattern indicates that an overwhelming proportion of evaluations have been restricted to short-range, homophilic tasks. Such a narrow focus risks overstating the general effectiveness of these methods and masking their potential limitations in long-range scenarios.
A few exceptions stand out. DropGNN [65], which was evaluated on graph-classification from the TUDataset [61], aligning more closely with evaluations of rewiring methods targeting over-squashing [44, 8]. NodeNorm, while primarily evaluated on homophilic datasets, was also tested on three heterophilic graphs: Cornell, Texas, and Wisconsin [66]. GCN-II saw broader evaluation, including on several long-range tasks such as Chameleon, Cornell, Texas, Wisconsin, and the Protein-Protein Interaction (PPI) networks [36]. Lastly, JKNet was also evaluated on the PPI networks.
A.3 Methods for Alleviating Over-squashing
@@
Spatial rewiring methods use the topological relationships between the nodes in order to come up with a rewiring strategy. That is the graph rewiring is guided by the objective of optimizing some chosen topological properties. For instance, Alon & Yahav [4] introduced a fully-adjacent (FA) layer, wherein messages are passed between all nodes. GNNs using a FA layer in the final message-passing step were shown to outperform the baselines on a variety of long-range tasks, revealing the importance of information exchange between far-off nodes which standard message-passing cannot facilitate. Topping et al. [79] proposed a curvature-based rewiring strategy, called the Stochastic Discrete Ricci Flow (SDRF), which aims to reduce the “bottleneckedness” of a graph by adding suitable edges, while simultaneously removing edges in an effort to preserve the statistical properties of the original topology. Black et al. [8] proposed the Greedy Total Resistance (GTR) technique, which optimizes the graph’s total resistance by greedily adding edges to achieve the greatest improvement. One concern with graph rewiring methods is that unmoderated densification of the graph, e.g. using a fully connected graph for propagating messages, can result in a loss of the inductive bias the topology provides, potentially leading to over-fitting. Accordingly, Gutteridge et al. [35] propose a Dynamically Rewired (DRew) message-passing framework that gradually densifies the graph. Specifically, in a given layer $\ell$, node $i$ aggregates messages from its entire $\ell$-hop receptive field instead of just the immediate neighbors. This results in an improved communication over long distances while also retaining the inductive bias of the shortest distance between nodes.
@@
Some other rewiring techniques don’t exactly classify as spatial or spectral methods. For instance, Probabilistically Rewired MPNN (PR-MPNN) [67] learns to probabilistically rewire a graph, effectively mitigating under-reaching as well as over-squashing. Finally, Gabrielsson et al. [25] proposed connecting all nodes at most $r$-hops away, for some $r\in\mathbb{N}$, and introducing positional embeddings to allow for distance-aware aggregation of messages.
A.4 Towards a Unified Treatment
@@
Several studies have shown that an inevitable trade-off exists between the problems of over-smoothing and over-squashing, meaning that optimizing for one will compromise the other. For instance, Topping et al. [79], Nguyen et al. [63] showed that negatively curved edges create bottlenecks in the graph resulting in over-squashing of information. On the other hand, Nguyen et al. [63, Proposition 4.3] showed that positively curved edges in a graph contribute towards the over-smoothing problem. To address this trade-off, they proposed Batch Ollivier-Ricci Flow (BORF), which adds new edges adjacent to the negatively curved ones, and simultaneously removes positively curved ones. In a similar vein, Giraldo et al. [34] demonstrated that the minimum number of message-passing steps required to reach a given level of over-smoothing is inversely related to the Cheeger’s constant, $h_{\mathcal{G}}$. This again implies an inverse relationship between over-smoothing and over-squashing. To effectively alleviate the two issues together, they proposed the Stochastic Jost and Liu Curvature Rewiring (SJLR) algorithm, which adds edges that result in high improvement in the curvature of existing edges, while simultaneously removing those that have low curvature.
@@
Despite the well-established trade-off between over-smoothing and over-squashing, some works have successfully tackled them together despite only adding or removing edges. One such work is Karhadkar et al. [44], which proposed a rewiring algorithm that adds edges to the graph but does not remove any. The First-order Spectral Rewiring (FoSR) algorithm computes, as the name suggests, a first-order approximation to the spectral gap of the symmetric Laplacian matrix ($\bm{L}^{\text{sym}}=\bm{I}_{N}-(\bm{D}^{\dagger})^{1/2}\bm{A}(\bm{D}^{\dagger})^{1/2}$), and adds edges with the aim of maximizing it. Since the spectral gap directly relates to Cheeger’s constant – a measure of bottleneck-edness in the graph – through Cheeger’s inequality [2, 3, 76], this directly decreases the over-squashing levels. Moreover, Karhadkar et al. [44, Figure 5] empirically demonstrated that addition of (up to a small number of) edges selected by FoSR can lower the Dirichlet energy of the representations, suggesting the method’s potential to simultaneously tackle over-smoothing. Taking a somewhat opposite approach, Liu et al. [54] adapted DropEdge to remove negatively curved edges sampled from a distribution proportional to edge curvatures. Their method, called CurvDrop, directly reduces over-squashing and, as a side benefit of operating on a sparser subgraph, also mitigates over-smoothing.
@@ Appendix B Proofs
B.1 Expected Propagation Matrix under DropEdge-variants
@@
Analysis of DropEdge-variants. We will similarly derive the expected propagation matrix for other random edge-dropping algorithms. First off, DropNode [23] samples nodes and drops corresponding columns from the aggregation matrix directly, followed by rescaling of its entries:
$\begin{split}\mathbb{E}_{\mathsf{DN}}\left[\frac{1}{1-q}\hat{\bm{A}}\right]=\frac{1}{1-q}\times\left(1-q\right)\hat{\bm{A}}=\hat{\bm{A}}\end{split}$ (B.15)
Nodes sampled by DropAgg [43] don’t aggregate messages. Therefore, if $\hat{\bm{A}}=\hat{\bm{A}}^{\text{asym}}$, then the expected propagation matrix is given by
$\begin{split}\mathbb{E}_{\mathsf{DA}}\left[\hat{\bm{A}}_{ii}\right]&=q+\frac{1-q}{d_{i}+1}=\frac{1+d_{i}q}{d_{i}+1}>\mathbb{E}_{\mathsf{DE}}\left[\hat{\bm{A}}_{ii}\right]\\ \mathbb{E}_{\mathsf{DA}}\left[\hat{\bm{A}}_{ij}\right]&=\frac{1}{d_{i}}\left(1-\mathbb{E}_{\mathsf{DA}}\left[\hat{\bm{A}}_{ii}\right]\right)<\mathbb{E}_{\mathsf{DE}}\left[\hat{\bm{A}}_{ij}\right]\end{split}$ (B.16)
Finally, DropGNN [65] samples nodes which neither propagate nor aggregate messages. From any node’s perspective, if it is not sampled, then its aggregation weights are computed as for DropEdge:
$\begin{split}\mathbb{E}_{\mathsf{DG}}\left[\hat{\bm{A}}_{ii}\right]&=q+\left(1-q\right)\mathbb{E}_{\mathsf{DE}}\left[\hat{\bm{A}}_{ii}\right]=q+\frac{1-q^{d_{i}+1}}{d_{i}+1}>\mathbb{E}_{\mathsf{DA}}\left[\hat{\bm{A}}_{ii}\right]\\ \mathbb{E}_{\mathsf{DG}}\left[\hat{\bm{A}}_{ij}\right]&=\frac{1}{d_{i}}\left(1-\mathbb{E}_{\mathsf{DG}}\left[\hat{\bm{A}}_{ii}\right]\right)<\mathbb{E}_{\mathsf{DA}}\left[\hat{\bm{A}}_{ij}\right]\end{split}$ (B.17)
@@ B.2 Sensitivity in L-layer Linear GCNs
where $\mathsf{ND}$ refers to a NoDrop model ($q=0$), the placeholder $\cdots$ can be replaced with one of the edge-dropping methods $\mathsf{DE}$, $\mathsf{DA}$ or $\mathsf{DG}$, and the corresponding entries of $\mathbb{E}_{\ldots}[\hat{\bm{A}}]$ can be plugged in from Equation 3.3, Equation B.16 and Equation B.17, respectively. Moreover, the sensitivity monotonically decreases as the dropping probability is increased.
@@
$\mathsf{Paths}\left(j\to i\right)=\left\{\left(u_{0},\ldots,u_{L}\right):u_{0}=j;u_{L}=i;\left(u_{\ell-1}\to u_{\ell}\right)\in\mathcal{E},\forall\ell\in\left[L\right]\right\}$ (B.19)
@@
The right hand side of the inequality is the $\left(i,j\right)$-entry in the Lth power of the propagation matrix of a NoDrop model. From Equation B.16 and Equation B.17, we know that Equation B.22 is true for DropAgg and DropGNN as well. We conclude the first part of the proof using Equation 3.2 – the sensitivity of node $i$ to node $j$ is proportional to $(\dot{\bm{P}}^{L})_{ij}$.
@@
Each of the terms on the right are increasing in $q$, hence, all the $\dot{\bm{P}}_{u_{\ell}u_{\ell-1}}$ factors are decreasing in $q$. Similarly, $\mathbb{E}_{\mathsf{DA}}[\hat{\bm{A}}_{ij}]$ and $\mathbb{E}_{\mathsf{DG}}[\hat{\bm{A}}_{ij}]$ decrease with increasing $q$. Using these results with Equation B.20, we conclude the second part of the theorem.
Appendix C Theoretical Extensions
C.1 Sensitivity in Nonlinear MPNNs
ReLU GCNs. Xu et al. [84] considered the case of ReLU nonlinearity, so that the update rule is $\bm{Z}^{\left(\ell\right)}=\mathsf{ReLU}(\hat{\bm{A}}\bm{Z}^{\left(\ell-1\right)}\bm{W}^{\left(\ell\right)})$. Additionally, it makes the simplifying assumption that each path in the computational graph is active with a fixed probability, $\rho$ [45, Assumption A1p-m]. Accordingly, the sensitivity (in expectation) between any two nodes is given as
@@ C.1 Sensitivity in Nonlinear MPNNs
$\left\|\mathbb{E}_{\mathsf{ReLU}}\left[\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right]\right\|_{1}=\left[\rho\left\|\prod_{\ell=1}^{L}\bm{W}^{\left(\ell\right)}\right\|_{1}\right]\left(\hat{\bm{A}}^{L}\right)_{ij}=\zeta_{1}^{\left(L\right)}\left(\hat{\bm{A}}^{L}\right)_{ij}$ (C.1)
where $\zeta_{1}^{\left(L\right)}$ depends only on the depth $L$, and is independent of the choice of nodes $i,j\in\mathcal{V}$. Taking an expectation w.r.t. the random edge masks, we get
@@
$\mathbb{E}_{\mathbf{M}^{\left(1\right)},\ldots,\mathbf{M}^{\left(L\right)}}\left[\left\|\mathbb{E}_{\mathsf{ReLU}}\left[\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right]\right\|_{1}\right]$ $=\zeta_{1}^{\left(L\right)}\left(\mathbb{E}\left[\prod_{\ell=1}^{L}\hat{\bm{A}}^{\left(\ell\right)}\right]\right)_{ij}=\zeta_{1}^{\left(L\right)}\left(\mathbb{E}\left[\hat{\bm{A}}\right]^{L}\right)_{ij}$ (C.2)
Using Theorem 3.1, we conclude that in a ReLU-GCN, DropEdge, DropAgg and DropGNN will reduce the expected sensitivity between nodes L-hops away. Empirical observations in s 1(a) and 4 suggest that we may expect an increase in sensitivity to neighboring nodes, but a significant decrease in sensitivity to those farther away.
Source-only Message Functions. Black et al. [8, Lemma 3.2] considers MPNNs with aggregation functions of the form
$\mathsf{Agg}^{\left(\ell\right)}\left(\bm{z}_{i}^{\left(\ell-1\right)},\left\{\bm{z}_{j}^{\left(\ell-1\right)}:j\in\mathbb{S}^{\left(1\right)}\left(i\right)\right\}\right)=\sum_{j\in\mathbb{B}^{\left(1\right)}\left(i\right)}\hat{\bm{A}}_{ij}\mathsf{Msg}^{\left(\ell\right)}\left(\bm{z}_{j}^{\left(\ell-1\right)}\right)$ (C.3)
$\left\|\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right\|_{1}\leq\zeta_{2}^{\left(L\right)}\left(\sum_{\ell=0}^{L}\hat{\bm{A}}^{\ell}\right)_{ij}$ (C.4)
$\mathbb{E}_{\mathbf{M}^{\left(1\right)},\ldots,\mathbf{M}^{\left(L\right)}}\left[\left\|\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right\|_{1}\right]$ $\leq\zeta_{2}^{\left(L\right)}\left(\mathbb{E}\left[I_{N}+\sum_{\ell=1}^{L}\prod_{k=1}^{\ell}\hat{\bm{A}}^{\left(k\right)}\right]\right)_{ij}=\zeta_{2}^{\left(L\right)}\left(\sum_{\ell=0}^{L}\mathbb{E}\left[\hat{\bm{A}}\right]^{\ell}\right)_{ij}$ (C.5)
Figure 5 shows the plot of the entries of $\sum_{\ell=0}^{6}\dot{\bm{P}}^{\ell}$ (i.e. for DropEdge), as in the upper bound above, with $\hat{\bm{A}}=\hat{\bm{A}}^{\text{asym}}$. We observe that the sensitivity between nearby nodes marginally increases, while that between distant nodes notably decreases (similar to 1(a)), suggesting significant over-squashing. Similar observations can be made with $\hat{\bm{A}}=\hat{\bm{A}}^{\text{sym}}$, and for other DropEdge-variants.
Source-and-Target Message Functions. Topping et al. [79, Lemma 1] showed that if the aggregation function is instead given by
$\mathsf{Agg}^{\left(\ell\right)}\left(\bm{z}_{i}^{\left(\ell-1\right)},\left\{\bm{z}_{j}^{\left(\ell-1\right)}:j\in\mathbb{S}^{\left(1\right)}\left(i\right)\right\}\right)=\sum_{j\in\mathbb{B}^{\left(1\right)}\left(i\right)}\hat{\bm{A}}_{ij}\mathsf{Msg}^{\left(\ell\right)}\left(\bm{z}_{i}^{\left(\ell-1\right)},\bm{z}_{j}^{\left(\ell-1\right)}\right)$ (C.6)
$\left\|\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right\|_{1}\leq\zeta_{3}^{\left(L\right)}\left(\hat{\bm{A}}^{L}\right)_{ij}$ (C.7)
With random edge-dropping, this bound can be adapted as follows:
$\mathbb{E}_{\mathbf{M}^{\left(1\right)},\ldots,\mathbf{M}^{\left(L\right)}}\left[\left\|\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right\|_{1}\right]\leq\zeta_{3}^{\left(L\right)}\left(\mathbb{E}\left[\hat{\bm{A}}\right]^{L}\right)_{ij}$ (C.8)
which is similar to Equation C.2, only with a different proportionality constant, that is anyway independent of the choice of nodes. Here, again, we invoke Theorem 3.1 to conclude that $(\mathbb{E}[\hat{\bm{A}}]^{L})_{ij}$ decreases monotonically with increasing DropEdge probability $q$. This implies that, in a non-linear MPNN with $\hat{\bm{A}}=\hat{\bm{A}}^{\text{asym}}$, DropEdge lowers the sensitivity bound given above. Empirical results in Figure 4 support the same conclusion for $\hat{\bm{A}}=\hat{\bm{A}}^{\text{sym}}$.
C.2 Test-time Monte-Carlo Dropout
Up until now, we have focused on the expected sensitivity of the stochastic representations in models using DropEdge-variants. This corresponds to their training-time behavior, wherein the activations are random. At test-time, the standard practice is to turn these methods off by setting $q=0$. However, this raises the over-smoothing levels back up [86]. Another way of making predictions is to perform multiple stochastic forward passes, as during training, and then averaging the model outputs. This is similar to Monte-Carlo Dropout, which is an efficient way of ensemble averaging in MLPs [27], CNNs [26] and RNNs [28]. In addition to alleviating over-smoothing, this approach also outperforms the standard implementation in practical settings [86]. We can study the effect of random edge-dropping in this setting by examining the sensitivity of the expected representations:
$\left\|\frac{\partial}{\partial\bm{x}_{j}}\mathbb{E}\left[\bm{z}^{\left(L\right)}_{i}\right]\right\|_{1}=\left\|\mathbb{E}\left[\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right]\right\|_{1}$ (C.9)
In linear models, the order of the two operations – expectation and 1-norm – is irrelevant:
$\left\|\mathbb{E}\left[\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right]\right\|_{1}=\left\|\mathbb{E}\left[\left(\hat{\bm{A}}^{L}\right)_{ij}\right]\bm{W}\right\|_{1}=\mathbb{E}\left[\left(\hat{\bm{A}}^{L}\right)_{ij}\left\|\bm{W}\right\|_{1}\right]=\mathbb{E}\left[\left\|\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right\|_{1}\right]$ (C.10)